# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

### v1

@@ 5.2 Main Results of SQL Generation (RQ1)
CodeS + KaSLA 68.32 50.75 38.19 60.17 64.52 96.37 89.46 76.44 63.86 84.82 84.46
StarCoder2 -15B TA-SL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
TA-SL + KaSLA 66.59 49.46 38.89 58.80 63.56 94.76 92.15 77.59 66.87 86.27 85.47
@@
DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 65.20 94.76 90.81 80.46 68.07 86.36 85.81
CodeS 68.00 51.40 39.58 60.30 65.04 94.76 91.26 72.99 64.46 84.72 83.52
CodeS + KaSLA 67.68 54.84 42.36 61.41 66.87 94.76 91.03 78.74 66.27 85.88 84.46
@@
E-SQL + KaSLA 69.40 53.55 43.75 62.18 65.32 95.97 92.38 73.56 65.06 85.69 85.72
@@
CHESS + KaSLA 69.94 54.84 47.92 63.30 67.21 96.37 92.83 74.14 65.66 86.17 86.35
Table 1: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets. We utilize KaSLA to substitute the original schema linking component in models that already incorporate schema linking, or to replace the full schema prompt for models that do not.
5.3 Experimental Results of Schema Linking (RQ2)
To answer the RQ2, we conduct the experiments and report the evaluation results of schema linking methods on BIRD-dev in Figure 2. We measured the performance of these schema linking methods in both table linking and column linking. The results indicate that, under the enhanced F1 score which specifically evaluates missing elements, only KaSLA achieves high accuracy in the challenging task of column linking. Schema linking models using LLMs often perform well in table linking but fall short in column linking due to the lack of additional optimizations for missing elements. This highlights the advantages of KaSLA’s framework and demonstrates how enhanced metrics can reveal the significant issue of element missing in current schema linking practices.
Schema linking method Table linking Column linking
DTS-SL (starcoder2-15b) 61.94 23.71
CodeS-SL (starcoder2-15b) 58.21 29.56
RSL-SQL (deepseek-v3) 67.11 17.20
KaSLA (deepseek-coder-1.3B) 81.92 62.78
Table 2: Schema linking performance with the proposed enhanced linking metric on BIRD-dev.
Ablation EX
CodeS + KaSLA 60.17
@@ 5.3 Experimental Results of Schema Linking (RQ2)
Table 3: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Schema linking model Binary scoring function with deepseek-coder-1.3B Probabilistic scoring model with RoBERTa-Large Estimation and dynamic programming Text-to-SQL Total time EX
BIRD-dev Full Schema / / / 5.15 s 5.15 s 57.63
KaSLA 3.5 s 0.12 s $<$ 0.01 s 1.85 s 5.48 s 60.17
Table 4: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
Model SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
BIRD-dev 58.60 60.30 60.17 60.02
@@
Table 5: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
5.4 Empirical studies (RQ3)
Ablation Study.
We presented the ablation study results for the key components in KaSLA in Table 3, the results concerning the choice of language model for the probabilistic scoring model are shown in Table 8 in Apendix and the scaling up performance of the scoring model are shown in Table 7 in Apendix, respectively.
As shown in Table 3: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones. This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified. (ii) Omitting the probabilistic scoring model risks missing potentially relevant elements, resulting in gaps within schema linking. The probabilistic model is essential for assigning a non-zero probability to all elements, thus capturing those with potential relevance. (iii) The hierarchical strategy effectively reduces the number of candidate columns, thereby lowering complexity and improving efficiency. By structuring the linking process in stages, KaSLA can more precisely target relevant schema components. (iv) Without the upper limit set at 1, the model introduces disproportionate comparisons among similarly relevant items, potentially excluding some relevant elements. This constraint is crucial to maintain balanced relevance assessments, preventing the overshadowing of items that might otherwise contribute positively to SQL generation. These ablation study highlights the significance of key components within KaSLA.
Transferability.
We conducted two experiments to evaluate KaSLA’s performance in cross-scenario transfer. We trained the binary and probabilistic scoring models on the Spider training dataset and evaluated it on the BIRD dev dataset, and vice versa, to explore its cross-scenario transfer ability. Results are provided in Table 5. We can find that pre-training KaSLA on public datasets yields results that outperform the baselines and are only slightly lower than what domain-specific fine-tuning would achieve. The results show that KaSLA demonstrates strong cross-scenario transferability, highlighting its robustness and adaptability across different data domains.
Inference time cost.
@@ 5.4 Empirical studies (RQ3)
We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table 4. By utilizing the lightweight LLM DeepSeek-coder-1.3B and significantly reducing the number of tokens in the input prompts through optimized schema linking, KaSLA achieves improved SQL generation accuracy with only a modest increase in inference time. This showcases KaSLA’s efficiency in balancing performance and computational demands.
6 Related Work

### v2

@@ V-A3 Implementation Details
CodeS + KaSLA 68.32 50.75 38.19 60.17 $\uparrow$ 4.41% 64.52 $\uparrow$ 2.06% 96.37 89.46 76.44 63.86 84.82 $\uparrow$ 0.12% 84.46 $\uparrow$ 1.13%
StarCoder2 -15B TA-SQL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
TA-SQL + KaSLA 66.59 49.46 38.89 58.80 $\uparrow$ 10.67% 63.56 $\uparrow$ 5.04% 94.76 92.15 77.59 66.87 86.27 $\uparrow$ 0.34% 85.47 $\uparrow$ 0.52%
@@
DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 $\uparrow$ 4.00% 65.20 $\uparrow$ 1.61% 94.76 90.81 80.46 68.07 86.36 $\uparrow$ 1.47% 85.81 $\uparrow$ 0.37%
@@
E-SQL + KaSLA 69.40 53.55 43.75 62.18 $\uparrow$ 3.46% 65.32 $\uparrow$ 5.08% 95.97 92.38 73.56 65.06 85.69 $\uparrow$ 0.56% 85.72 $\uparrow$ 2.28%
@@
CHESS + KaSLA 69.94 54.84 47.92 63.30 $\uparrow$ 3.20% 67.21 $\uparrow$ 2.96% 96.37 92.83 74.14 65.66 86.17 $\uparrow$ 0.78% 86.35 $\uparrow$ 1.20%
RSL-SQL 69.95 54.62 47.22 63.17 67.11 96.37 93.05 74.71 66.87 86.56 86.42
RSL-SQL + KaSLA 71.03 56.34 51.39 64.73 $\uparrow$ 2.47% 68.12 $\uparrow$ 1.35% 97.18 93.50 75.29 68.07 87.14 $\uparrow$ 1.50% 87.36 $\uparrow$ 1.09%
V-B Main Results of SQL Generation (RQ1)
To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table I. KaSLA serves as a plug-in method: for models with built-in schema linking modules, it replaces their original linking component; for models without a schema linking component, it substitutes the conventional full schema prompting approach.
TABLE II: Schema linking performance with Recall+, Precision+, and F1+ on BIRD-dev. We extracted the schema linking components from the baselines and used “-SL” instead of the original “-SQL” to indicate this.
Dataset Metric Recall-based Method Generative Method KaSLA
CodeS-SL ICL DIN-SL DTS-SL CHESS-SL E-SL Pure-SL TA-SL RSL-SL
BIRD-dev Recall+ Table 99.15 88.40 82.66 64.99 80.70 69.04 64.99 79.47 92.57 94.52
Column 85.14 22.03 16.17 62.52 14.93 22.36 41.33 54.69 57.11 81.29
Precision+ Table 43.03 82.75 78.80 60.29 49.92 68.01 60.29 74.49 83.89 75.00
Column 18.65 20.69 15.12 16.50 7.09 21.57 37.23 52.31 35.14 53.53
F1+ Table 58.21 84.79 80.18 61.94 58.51 68.39 61.94 76.16 86.93 81.92
Column 29.56 21.20 15.51 23.71 9.03 21.87 38.77 53.20 40.60 62.78
The results in Table I demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3. Notably, on the BIRD-dev dataset, KaSLA raises the EX of RSL-SQL with Deepseek-v3 from 63.17% to 64.73%, achieving a 2.47% improvement and the best overall result. KaSLA also delivers a 3.20% increase on CHESS. For local LLMs, both “CodeS + KaSLA” with CodeS-15B and “TA-SL + KaSLA” with StarCoder-15B achieve significant gains in Execution Accuracy compared to their original versions. We also found that the differences between baselines on Spider-dev are smaller compared to BIRD-dev. This may be because the schemas in Spider-dev are less complex. Even in this case, KaSLA can still consistently provide improvements. These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks. This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing text-to-SQL systems.
V-C Experimental Results of Schema Linking (RQ2)
To answer the RQ2, we conduct the evaluation of schema linking methods on BIRD-dev using the proposed schema linking metrics. Table II shows the performance of various schema linking methods on BIRD-dev, evaluated by Recall+, Precision+ and F1+ at both the table and column levels. The results allow us to compare the schema linking performance of recall-based, generative, and the proposed KaSLA method.
As shown in Table II, most current schema linking models perform well in table linking, but linking relevant columns from a large number of candidate columns remains challenging. The recall-based method (CodeS-SL) achieves the highest Recall+ at both the table and column levels, indicating it is very effective at identifying relevant elements and avoiding missing important ones. However, this comes at the cost of low Precision+, especially for columns, meaning many redundant elements are also linked. As a result, its F1+ scores for both tables and columns are low. Generative methods tend to optimize for Precision+, but their Recall+ is much lower than that of the recall-based method, indicating more missing elements. Even the best-performing generative model, RSL-SL, achieves a low Precision+ and, therefore, a low F1+ on the challenging column linking task. Notably, models like E-SQL and CHESS, which achieve strong results in Table I, obtain low scores for column linking when their schema linking modules are evaluated separately. This suggests that E-SQL uses LLMs to process the full schema directly without schema linking [26] and CHESS reduces error rates through multiple rounds of iteration [11], which both result in high LLM API costs. KaSLA demonstrates strong and balanced performance. Compared to generative models, it achieves higher Recall+, only slightly lower than the recall-based method, while also obtaining competitive Precision+ and the best F1+ at the column level, as well as a strong table-level F1+. These results show that KaSLA achieves the best overall trade-off by effectively balancing recall and precision, successfully reducing missing and redundant elements simultaneously. This allows KaSLA to outperform recall-based and generative baselines in schema linking quality, particularly for the challenging task of column linking, thus demonstrating the advantage of its knapsack-optimization framework for text-to-SQL tasks.
V-D Ablation Study (RQ3)
To answer RQ3, we present: (i) the ablation study results for KaSLA’s key components, (ii) the performance analysis with different language model choices for the probabilistic scoring model, and (iii) the effects of scaling up the scoring model.
V-D1 Key Components in KaSLA
As shown in Table III: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones. This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified. (ii) Omitting the probabilistic scoring model risks missing relevant elements, because the soft relevance probabilities it provides can effectively protect against the mistakes made by the binary scoring model with hard relevance scores. It provides a more robust estimation. (iii) The hierarchical strategy effectively reduces the number of candidate columns. Without this strategy, column linking would have to be performed on all candidate columns, which introduces extra noise and increases computational cost. (iv) Without the upper limit set at 1, the model may introduce disproportionate comparisons among relevant items, potentially excluding some relevant ones. This constraint is crucial to maintain balanced relevance estimation, preventing the overshadowing of elements that might otherwise contribute positively to SQL generation. These ablation study highlights the significance of key components within KaSLA.
TABLE III: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS-15B as the text-to-SQL backbone.
Dataset Ablation of key components of KaSLA EX
BIRD-dev CodeS + KaSLA 60.17
@@ V-D1 Key Components in KaSLA
(a) Binary scoring model
(b) Probabilistic scoring model
Figure 6: (a) Execution accuracy (EX, %) of CodeS + KaSLA on BIRD-dev with different language models as the binary scoring model. (b) Effect of model parameter scaling of probabilistic scoring model with Deepseek-coder. All experiments use CodeS-15B as the text-to-SQL backbone.
V-D2 Binary Scoring Model
We present ablation studies on the choice of language models for the binary scoring model used in KaSLA, as detailed in Figure 6(a). We evaluate the SQL generation performance of CodeS-15B on BIRD-dev using KaSLA with different language models as the binary scoring model. In Figure 6(a), we find that RoBERTa-Large achieves the best results. One reason is that RoBERTa enhances BERT by employing a more comprehensive training process, leading to improved text representation abilities [16]. This allows RoBERTa-Large to provide better generalization and understanding of language nuances, which are crucial for schema linking tasks where capturing subtle semantic relationships is important. Moreover, reducing the number of parameters (e.g., using BERT-base or RoBERTa-base) resulted in relatively underfitted models that failed to capture sufficient semantic information for the task. Conversely, increasing the number of parameters (e.g., SGPT-1.3B) introduced a tendency toward overfitting, which did not yield performance improvements. These results indicate that RoBERTa-Large offers an effective balance between model capacity and generalization, making it an excellent choice for the binary scoring model.
TABLE IV: The comparison of schema linking performance between CodeS-SL and KaSLA is shown using a randomly selected case from BIRD-dev. We highlight the missing relevant element in CodeS-SL’s schema linking results in blue, and the redundant elements that mislead the SQL generation model in red. These two types of errors together result in the incorrect SQL generation of CodeS-SL. In contrast, KaSLA successfully avoids these issues.
Question ‘Posters’ is the expense description; ‘September Speaker’ is an event name; Indicate the cost of posters for ‘September Speaker’ event.
Method CodeS-SL KaSLA
Schema linking result event: [event_id, event_name, event_date, status, notes, location] zip_code: [type, county, city, state, zip_code, short_state] expense: [cost, expense_description, expense_id, expense_date, approved, link_to_budget] attendance: [link_to_member, link_to_event] budget: [event_status, category, spent, amount, remaining, link_to_event] event: [event_id, event_name] expense: [cost, expense_description, expense_date, link_to_budget] budget: [budget_id, link_to_event]
Missing elements budget: [budget_id] \
Redundant elements event: [event_date, status, notes, location] zip_code: [type, county, city, state, zip_code, short_state] expense: [expense_id, expense_date, approved] attendance: [link_to_member, link_to_event] budget: [event_status, category, spent, amount, remaining] expense: [expense_date]
Generated SQL using CodeS-15B SELECT expense.cost FROM expense INNER JOIN attendance ON expense.link_to_member = attendance.link_to_member INNER JOIN event ON attendance.link_to_event = event.event_id WHERE event.event_name = ‘September Speaker’ AND expense.expense_description = ‘Posters’ SELECT expense.cost FROM event INNER JOIN budget ON event.event_id = budget.link_to_event INNER JOIN expense ON budget.budget_id = expense.link_to_budget WHERE event.event_name = ‘September Speaker’ AND expense.expense_description = ‘Posters’
SQL execution result Incorrect Correct
V-D3 Probabilistic Scoring Model
We present ablation studies on LLM selection with parameter scaling for KaSLA’s probabilistic scoring component in Figure 6(b). We find that although using larger LLMs as the probabilistic scoring model leads to improved performance, the difference in execution accuracy between Deepseek-coder-1.3B and the larger 6.5B and 7B models is marginal. Importantly, Deepseek-coder-1.3B offers clear advantages in terms of lower resource consumption and faster inference time due to its smaller parameter size. These results indicate that, within the knapsack optimization framework of KaSLA, a low-parameter LLM can achieve performance comparable to that of much larger models. This finding demonstrates that KaSLA is well-suited for efficient and scalable deployment in practical text-to-SQL applications, where smaller models are often preferred due to constraints on computational resources and the need for faster inference.
V-E Empirical Studies (RQ4)
To address RQ4, we conduct a series of empirical studies to evaluate the practical effectiveness of KaSLA. First, we provide a randomly selected case study in Table IV to demonstrate how KaSLA improves SQL generation in real scenarios compared with a baseline. Next, we evaluate KaSLA’s transferability across domains and scenarios, with the results summarized in Table V and Table VI, respectively. The results show that KaSLA has strong transferability on unseen data. We also offer a complexity analysis and present the inference time cost when integrating KaSLA into text-to-SQL frameworks in Table VII, highlighting KaSLA’s ability to enhance SQL generation accuracy with only a modest increase in inference time. Finally, we systematically analyze the failure cases of “CodeS + KaSLA” on the BIRD-dev, as illustrated in Figure 8, to provide a comprehensive understanding of KaSLA’s schema linking performance in SQL generation tasks.
V-E1 Case Study
To illustrate KaSLA’s practical impact, we provide a randomly selected case study ib BIRD-dev dataset to illustrate how KaSLA improves SQL generation in practice, comparing with CodeS-SL. The case are detailed in Table IV, it demonstrates how KaSLA effectively avoids missing relevant elements compared to the baseline. The original schema linking model in CodeS introduced many redundant elements while missing the relevant column $budget\_id$ in the $budget$ table. After replacing it with KaSLA, the schema linking results included all relevant elements with minimal redundancy, enabling the SQL generation model to produce the correct query. This case highlights KaSLA’s ability to handle complex, real-world scenarios, which complements the quantitative results with tangible evidence of its utility.
Figure 7: The seen and unseen data in BIRD-dev relative to BIRD-train.
TABLE V: Execution Accuracy (EX) (%) of text-to-SQL models evaluated on seen data and un-seen data in BIRD-dev.
Evaluation dataset SuperSQL CodeS CodeS + KaSLA
Original BIRD-dev 58.60 57.63 60.17
Seen data in BIRD-dev 69.90 67.58 71.06
Un-seen data in BIRD-dev 49.35 51.01 86.48
V-E2 Transferability
In real-world scenarios, databases come from diverse domains, leading to varied data and queries. Unlike open-source datasets, real-world tasks often lack enough in-domain training data, so cross-domain transferability of schema linking models is essential. We specifically evaluate KaSLA’s performance in cross-domain and cross-scenario settings.
Cross-domain Transferability
The BIRD dataset is specifically designed for cross-domain evaluation. As shown in Figure 7, there are 5 databases in BIRD-dev whose domains match those in BIRD-train. For example, “european_football_2” in BIRD-dev matches with “european_football_1” in BIRD-train. More relations between data from BIRD-dev and BIRD-train are illustrated in Figure 7. This type of data is considered seen data, and the model can learn prior knowledge for these domains from BIRD-train. In contrast, there are six databases in BIRD-dev that do not have directly related domains in BIRD-train. For example, “card_games” in BIRD-dev does not have any domain-matched database in BIRD-train. For these databases, the model cannot acquire relevant prior knowledge through training. Evaluating KaSLA on these unseen domains provides a actual test of its cross-domain transferability.
We evaluated the SQL generation performance of KaSLA and baseline models on data from both the unseen and seen databases in BIRD-dev. As shown in Table V, KaSLA demonstrates a clear advantage over baselines on both seen and unseen data. This strong performance is attributed to the robustness of the KaSLA optimization framework, which uses knapsack optimization to reduce missing and redundant elements. It enables KaSLA to achieve optimal results even on unseen domains after training on open-source datasets.
Cross-scenario Transferability
Given the diversity of databases in real-world scenarios, evaluating KaSLA’s performance under cross-scenario conditions provides a more practical assessment of its potential for real-world text-to-SQL applications. Since BIRD and Spider are datasets from different scenarios, we assess KaSLA’s cross-scenario transferability by evaluating KaSLA trained on BIRD-train with Spider-dev, and vice versa—evaluating KaSLA trained on Spider-train with BIRD-dev. For baselines, we follow the standard setting of using training or prompt sampling from the same scenario as the evaluation data. Results are shown in Table VI. Here, “KaSLA (Transfer)” denotes KaSLA trained on the training set of a scenario different from the evaluation data.
We observe that KaSLA trained on a different-scenario dataset still outperforms the baselines and only slightly underperforms compared to KaSLA trained with domain-specific data from the matching scenario. These results demonstrate that KaSLA has strong cross-scenario transferability, highlighting its robustness and adaptability when applied to new text-to-SQL scenarios. This also confirms that KaSLA, when trained on open-source datasets, can be directly and effectively applied to unseen scenarios in real-world applications.
TABLE VI: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
Evaluation dataset SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
BIRD-dev 58.60 57.63 60.17 60.02
@@ V-E2 Transferability
V-E3 Complexity Analysis and Inference Time Cost
We provide a complexity analysis for offering insights into KaSLA’s implementation and computational efficiency. The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism. KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables. For a database with $|{\mathcal{T}}|$ tables and a table tolerance of $u^{{\mathcal{T}}}$, the table linking phase has a time complexity of $O(|{\mathcal{T}}|u^{{\mathcal{T}}})$. KaSLA then links columns within the selected table. Let $|{\mathcal{C}}^{t}|$ and $u^{{\mathcal{C}}}$ represent the number of columns and the column capacity of a selected table, respectively. The time complexity of the column linking phase is $O(|{\mathcal{C}}^{t}|u^{{\mathcal{C}}})$. Consequently, the total time complexity of KaSLA is $O(|{\mathcal{T}}|u^{{\mathcal{T}}}+|{\mathcal{T}}|(|{\mathcal{C}}^{t}|u^{{\mathcal{C}}}))$. Since all factors are constants, this results in linear complexity.
@@ V-E3 Complexity Analysis and Inference Time Cost
We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table VII. By utilizing the lightweight LLM DeepSeek-coder-1.3B and significantly reducing the number of tokens in the input prompts through optimized schema linking, KaSLA achieves improved SQL generation accuracy with only a modest increase in inference time. This showcases KaSLA’s efficiency in balancing performance and computational demands.
TABLE VII: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
Model Full Schema KaSLA
Binary scoring function / 3.5 s
Probabilistic scoring model / 0.12 s
Estimation and dynamic programming / $<$ 0.01 s
SQL generation with CodeS-15B 5.15 s 1.85 s
Total time 5.15 s 5.48 s
EX 57.63 60.17
V-E4 Failure Case Analysis
Figure 8: Failure case analysis of “CodeS + KaSLA” in BRD-dev.
We present a systematic analysis of the failure cases of “CodeS + KaSLA” in BRD-dev, as detailed in Figure 8, to comprehensively study the schema linking performance of KaSLA. We first categorize all evaluation results into two groups: correct SQL generation and incorrect SQL generation. Next, we further divide the cases of incorrect SQL generation into three subcategories: (a) schema linking performs well, but the SQL generation model is insufficient; (b) the SQL generation model is misled by redundant elements; and (c) the SQL generation model is misled by missing elements. Notably, subcategories (b) and (c) are directly caused by errors in the schema linking model, which subsequently lead to incorrect SQL generation. In contrast, in category (a), the schema linking provides correct linking results, yet the SQL generation model still produces incorrect SQL.
As show in Figure 8, in 79.40 % of examples (60.17% + 19.23%), KaSLA provides schema links without missing or misleading redundant elements. However, while 60.17% of these cases yield correct SQL, 19.23% still result in incorrect SQL due to limitations in the SQL generation model. This suggests that adopting more advanced SQL generation models could further improve the text-to-SQL performance. We can also observe that, despite KaSLA’s tailored optimization for schema linking, it still misses relevant schema elements in 9.52% of examples and introduces redundant elements, which mislead the SQL generation model in 11.08% of cases. This indicates room for improvement in the future work, particularly in refining the knapsack optimization algorithm or enhancing the relevance scoring model’s accuracy to reduce redundancies while minimizing the omission of relevant elements.
VI Related Work
VI-A LLM-based Text-to-SQL
Alongside the emergence of LLMs, in-context learning (ICL), also known as prompt engineering has become a core method driving advancements [29]. Early efforts in LLM-based text-to-SQL studied the effectiveness of ICL with different prompt designs on various LLMs’ structures [30]. With the success of well-designed ICL methods in other natural language understanding tasks [31, 32], their integration into text-to-SQL has also achieved solid improvements [33]. Chain-of-Thought (CoT) prompting [31] shows great potential for text-to-SQL tasks: DIN-SQL [22] and ACT-SQL [34] proposed a CoT-based framework to enhance SQL generation. To optimize the prompt, DAIL-SQL [4] and Knowledge-to-SQL [35] provide related samples and helpful knowledge for LLMs to learn with. Furthermore, MAC-SQL [23] and CHESS [11] employs a multi-agent collaboration framework to generate and refine SQL; CHASE-SQL [36] proposes an ICL prompt for question decomposition and involves a novel online synthetic example generation. ICL-based methods are the mainstream paradigm of LLM-based text-to-SQL, which have made significant progress and continue to be widely studied in the most current work [3].
As another paradigm, fine-tuning is an intuitive and widely recognized technique for enabling LLMs to perform specific downstream tasks [32]. Even though code-specific LLMs are trained on extensive programming scenarios, they still struggle when facing challenging user queries and complex database environments [37, 38, 19]. As a straightforward method, supervised fine-tuning (SFT) [32] is utilized to adapt open-source LLMs for SQL generation [4]. For well-designed methods, CodeS [8] proposes a two-stage training framework achieving impressive performance with open-source LLMs; PURPLE [39] retrieves demonstrations containing the requisite logical operator composition to guide LLMs for SQL generation. DTS-SQL [9] fine-tunes two LLMs separately for schema linking generation and SQL generation; XiYan-SQL [40] proposes a multi-generator ensemble framework through fine-tuning to enhance SQL candidate quality and diversity, with a refined selection model and schema representation.
VI-B Schema Linking in Text-to-SQL